an ideal heat engine has an efficiency of 25% when its low temperature reserver is 25°C. it is desire to increase the efficiency to 35%. what change should be made in either (1) the temperature of the low reserver or (2) the temperature of the high reserver??

1
Expert's answer
2021-11-30T09:49:52-0500

The efficiency is


"\eta=1-\frac{T_c}{T_h},\\\space\\
T_h=\frac{T_c}{1-\eta}=397\text{ K}"

for 25%. For 35%, the following changes can be made:

"T_h=\frac{T_c}{1-\eta}=458\text{ K},\\\space\\
T_c=(1-\eta)T_h=258\text{ K}"

to the temperatures of the hot and cold receiver respectively.